OpenStudy (anonymous):

x-1/x as x approaches 0

OpenStudy (owlcoffee):

\[\lim_{x \rightarrow 0}\frac{ x-1 }{x} = \] knowing that we can't divide by 0, the limit would be an indetermination.

OpenStudy (anonymous):

But wouldnt it be infinite since o approaches 0

OpenStudy (anonymous):

Wrong, this is not an indeterminate form: \[ \frac {-1}0 \]We know that this is undefined!

OpenStudy (anonymous):

Consider \[ x\times 0=-1 \]It is not like \[ x\times 0=0 \]

OpenStudy (anonymous):

In the first equation, there is NO defined number that could be a solution for \(x\). Hence undefined. In the second equation there a many possible solutions for \(x\) and we don't have the context to determine which is correct. Hence indeterminate.

OpenStudy (anonymous):

This limit does not exist. It can approach \(-\infty\) or \(\infty\) depending which side you come from.

OpenStudy (abb0t):

Nicely said @wio

OpenStudy (anonymous):

oh got it but the beginning was kind of unclear but thanks